# Gas contamination and mitigation in a 100 m3 / 10 bar argon TPC with optical readout: a viability study

### v1

@@ 3.2 CFD simulations
A numerical evaluation, based on the well established CFD methodology, allows for a realistic approach to describe the gas and impurity dynamics by calculating the velocity field in the overall volume of the field cage, including the injection setup. It is assumed that impurities, present in trace amount, do not modify the velocity field of the main gas. Translational invariance along the (axial) $z$ dimension $(\frac{\partial}{\partial z}=0)$ was assumed. This choice clings on the (axial) orientation of the distributors and the discharge optimization performed in the previous section, that assures the discharge angle to be nearly perpendicular to the distributor (table 2). As discussed, the outgassing rate from the end-caps can be made sub-dominant relative to that from the field cage, approximately preserving the adopted symmetry. Although the region between holes cannot be well described in this approach, the distance chosen in geometries $A$-$D$ in table 2 is sufficiently small (15-50 mm) to anticipate a negligible (local, at most) impact.
@@
The hole aperture of the inlets and local pressure difference (distributor-TPC) were set to provide (simultaneously) the desired average velocity per hole and flow per unit length, based on the values in table 2. Achieving convergence in the simulations by setting the hole velocity instead of the pressure difference as an initial condition was not possible in general. The outlets were set to a fixed pressure of 10 bar. Gravity was included in simulation, and in particular the hydrostatic pressure was added automatically at the inlets and outlets depending on their height. Outgassing was set perpendicular to the walls and to the distributors. As shown in Fig. 4, the mesh is finer closer to the holes and coarser towards the center of the field cage. To assure convergence, the solution was checked to be stable upon global up and down-scalings on the mesh size and variations on the time step.
@@
The turbulent-flow ‘k-$\varepsilon$’ model was selected, being adequate for simulating single-phase flows at high Reynolds numbers and suitable for incompressible or weakly compressible flows, and compressible flows at low Mach numbers. The ‘k-$\varepsilon$’ model is the most validated turbulence model and performs well in confined flows where the Reynolds shear stresses are important. This model uses the Reynolds-averaged Navier-Stokes equations (RANS), introducing two more magnitudes: the turbulent energy ($k$) and the rate of turbulent energy loss ($\varepsilon$). These magnitudes have their corresponding differential equations coupling the velocity field ($\overrightarrow{v}$) with the viscous stress tensor K (following COMSOL notation):
$\rho\overrightarrow{\nabla}\cdot\overrightarrow{v}=0,$ (18)
$\rho\frac{\partial\overrightarrow{v}}{\partial t}+\rho(\overrightarrow{v}\cdot\overrightarrow{\nabla})\overrightarrow{v}=\overrightarrow{\nabla}\cdot\left[-P\textbf{I}+\textbf{K}\right]+\rho\overrightarrow{g},$ (19)
$\rho\frac{\partial k}{\partial t}+\rho(\overrightarrow{v}\cdot\overrightarrow{\nabla})k=\overrightarrow{\nabla}\cdot\left[\left(\mu+\frac{\mu_{T}}{\sigma_{k}}\right)+\overrightarrow{\nabla}k\right]+P_{k}-\rho\varepsilon,$ (20)
$\rho\frac{\partial\varepsilon}{\partial t}\!+\!\rho(\overrightarrow{v}\!\cdot\!\overrightarrow{\nabla})\varepsilon=\overrightarrow{\nabla}\!\cdot\!\left[\left(\mu+\frac{\mu_{T}}{\sigma_{\varepsilon}}\right)\!+\overrightarrow{\nabla}\varepsilon\right]+C_{\varepsilon 1}\frac{\varepsilon}{k}P_{k}-C_{\varepsilon 2}\rho\frac{\varepsilon^{2}}{k}.$ (21)
In the above expressions, 18 is the continuity equation for incompressible flow, I is the identity matrix and $\overrightarrow{g}$ is the gravity force. The following definitions are needed:
$\mu_{T}=\rho C_{\mu}\frac{k^{2}}{\varepsilon};~{}~{}~{}P_{k}=\mu_{T}\left[\overrightarrow{\nabla}\overrightarrow{v}:(\overrightarrow{\nabla}\overrightarrow{v}+(\overrightarrow{\nabla}\overrightarrow{v})^{T})\right],$ (22)
where $\mu_{T}$ is the turbulent viscosity (a correction to the viscosity $\mu$ because of the turbulent regime), $P$ its pressure, and the inner product is defined as:
$\textbf{a}:\textbf{b}=\sum_{i}\sum_{j}a_{ij}b_{ij}.$ (23)
The dimensionless coefficients, called closure coefficients, of these equations are [47]:
@@
$\overrightarrow{J}=-D\overrightarrow{\nabla}c,$ (24)
@@
$\frac{\partial c}{\partial t}+\overrightarrow{\nabla}\cdot\overrightarrow{J}+\overrightarrow{v}\cdot\overrightarrow{\nabla}c=0,$ (25)
and diffusion coefficients are taken from the last column in table 1. At the surface boundaries, the ‘no slip’ condition was used [50], as it is customary for solid boundaries and Newtonian fluids in macro-scale flow and non-rarefied conditions.
@@ 4.1 TPC filling
A priori, the displacement of air during the chamber filling can be performed at flow rates higher than those imposed by the recirculation compressors, if need be. In the ideal case, it might even be possible to pump the TPC, although for large volumes this cannot be taken for granted. Here we consider for illustration the process of TPC filling at a flow rate of 6.5 m3/h (‘typical’ flow), in the default geometry ($B$). It has been assumed, for simplicity, that the chamber is filled quickly (and uniformly) up to 9 bar of Ar and 1 bar of air (by making use of the large pressure difference at the start of injection) and afterwards the flow of argon is set to the nominal one (that marks the time $t=0$ h in simulation). In the absence of convection, the triple-hole distributor proposed in geometry $B$ removes very efficiently the O2 concentration from the region in between distributors and then the fresh argon gently displaces the Ar/O2 admixture upwards. The relatively low diffusion of O2 at high pressure and its lower mass favors its extraction from the uppermost outlets.
Figure 8: Evolution of the oxygen concentration upon starting the gas flow in the TPC ($\Phi=6.5$ m3/h, $P=10$ bar), for distributor geometry $B$. The initial value corresponds to a 21% O2 concentration and 1 bar of air / 9 bar of argon. The line shows the 0.1 ppm landmark, corresponding to a 10% charge loss in a cathode-anode electron transit.
Fig. 8 shows the concentration of oxygen integrated over the chamber as a function of time. It is worth noting that, in the limit of large diffusion (where O2 would be uniformly distributed over the chamber at all times), the time constant for O2 replacement would be directly $\tau_{{}_{\text{O}_{2}}}=V/\Phi$ (see appendix), with $V$ being the chamber volume (100 m3). Reaching the 0.1 ppm landmark value, for instance, would take 188 h, about ×$\times 4.5$ more time than observed in the CFD simulations (42 h). While it is clear that convection currents due to power dissipation by active elements can slow down the displacement of oxygen in the final experimental conditions, it is safe to state that, in the absence of pumping, the filling time will be situated in the range 40-190 h. An alternative based on a modest increase of the gas flow during filling could be viable; however, a significant volume of gas will need to be disposed of, purified or recovered in real time.
Figure 9: Time lapse of the distribution of O2 concentration in our reference 10 bar Ar/CF4 TPC for geometry $B$ (holes in ‘triplet’ configuration), under typical outgassing and flow rate conditions. The concentric circles mark the limits of the different mesh regions. The transient phase of the gas mixing can be observed mostly during the first 20 h. After that, the system approaches steady state and the concentration becomes homogeneous in the chamber. The images show clearly how the gas mixing is driven in these conditions by the velocity field, going from bottom to top. (The color bar displays the logarithm of the concentration in mol/m3.)
@@ 4.2 Oxygen contamination
When studying the evolution of oxygen, two outgassing rates have been considered: i) $Q=10^{-8}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$ (‘typical’) and ii) $Q=0.05\times 10^{-8}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$ (‘optimistic’). Similarly, for the gas flow: i) $\Phi=6.5$ m3/h (‘typical’) and ii) $\Phi=20$ m3/h (‘optimistic’). The outgassing rate was assumed constant and equal to its value at 10 h. This has a dual benefit: i) it provides a safe upper limit to the O2 concentration (compared to the anticipated $t^{-1/2}$ law) and ii) it allows a straightforward interpretation of the moment when the system achieves stationary conditions. Details are compiled in Table 1.
Figure 10: Time lapse of the distribution of O2 concentration in our reference 10 bar Ar/CF4 TPC for geometry $A$ (radial holes), under typical outgassing and flow rate conditions. After 24 h, steady state conditions have not been achieved yet. (The color bar displays the logarithm of the concentration in mol/m3.
@@
Fig. 9 shows the evolution of the O2 concentration under typical outgassing and flow rate conditions for geometry $B$ (triple-hole distributor). For comparison, Fig. 10 shows the results for geometry $A$ (single-hole distributor) when the holes are radially-oriented. This latter case displays the formation of a vortex. Therefore, while in the default geometry $B$ (hole triplet) the contaminants are uniformly mixed quickly thanks to the velocity field, in geometry $A$ (hole singlet) the uniformity is achieved through diffusion, involving considerably longer times. Single or double vortexes are prevalent in single-hole distributor geometries, for the conditions studied, but are absent in the triple-hole ones. The implications of this observation can be summarized in Fig. 11. In the left panel, the time evolution of the O2 concentration is shown for three different distributor geometries (triple holes: blue; radial holes: orange; lined-up holes: green), corresponding to geometries $A$ and $B$ in table 2. The right panel shows the radial distribution of impurities (logarithmic X-scale) when all geometries have achieved steady state (500 h): it is completely homogeneous except in the outermost regions, where it can vary up to 30%. The triplet configuration reaches stationary conditions within 20 h, compared to almost ×$\times 10$ that time for the 1-hole geometries. [^3] The stationary limit following eq. 43 in appendix is $\frac{Q\cdot{A}}{P\cdot\Phi}=0.43$ ppm, to be compared with the simulated one of 0.31 ppm. The deviation can be attributed to i) the uncertainty in the flow estimate from the average velocity at the inlets and ii) the contribution from the last 5 cm of the chamber, that has a complex dynamics and has been removed here aiming at a simplified discussion (the last cm’s from the chamber are usually excluded from any physics analysis).
@@
It is difficult to assess the significance of the larger tendency to vortex formation in the single-hole distributors, and whether it will be maintained in an actual experiment when considering the exact geometry and the unavoidable convection currents stemming from temperature gradients. Notwithstanding, the triple-hole distributor is vortex-free and achieves consistently a faster mixing with contaminants as well as a more efficient sweeping of the outgassing from the chamber walls. This is illustrated in Fig. 12, where three different gas flows are compared under typical values of the outgassing rate (top figure: 2 m3/h, middle figure: 6.5 m3/h, bottom figure: 20 m3/h). The time evolution is displayed, up to 48 h, on the left column, while the radial distribution of contaminants is shown on the right column. Due to the low gas flow in the top figure (and correspondingly low discharge velocity), the time scale is too short to lead to vortex formation and the two distribution schemes perform similarly up to nearly the steady state value (1.4 ppm for these conditions). Even in the absence of vortex, the triple-hole distribution is more efficient at mixing due to a better sweeping of the chamber walls. For larger gas flows (middle and bottom figures), the fluid dynamics of the radial-hole geometry is dominated by the formation of a central vortex, with impurities reaching the central region through diffusion alone (right figures).
Figure 13: Time evolution of the O2 contamination in our reference 10 bar Ar/CF4 TPC for ‘typical’ outgassing and flow rates, in the default geometry $B$ (holes in ‘triplet’ configuration) in red. A 1-D solution (‘infinite diffusion’ limit) is shown in dashed lines. The CFD solution for ‘optimistic’ outgassing values (blue) reflects a simple down-scaling, without altering the shape. (To account for the numerical deviation from the steady-state limit, as discuss earlier in text, the outgassing in the 1-D solution has been downscaled by 28%.)
The results obtained in this section can be analyzed by comparison with the one-dimensional solutions, that effectively consider an ‘infinite diffusion’ scenario (see appendix). Fig. 13 shows how the observed time-dependence (red line) is reasonably well approximated by the 1-D solution (dashed line), implying that the achieved mixing rate is compatible with the ideal limit of perfect mixing. The blue line, obtained at ×$\times 20$ less outgassing (‘optimistic’ case) aims at illustrating that the curves obtained by CFD reflect a global scaling factor for different outgassing rates, following as well the expectation from the 1-D solutions (see appendix). The asymptotic value in Fig. 13 (red line) of 0.3 ppm would correspond to a 30% charge loss following the estimates in section 2.1. Considering the range of outgassing rates in table 1, an upper limit to the charge loss after 24 h can be situated in the range 1.5-30%, for the gas flow conditions in Fig. 2.1.

### v2

@@ 3.2 CFD simulations
A numerical evaluation, based on the well-established CFD methodology, allows for a realistic approach to describe the gas and impurity dynamics by calculating the velocity field in the overall volume of the field cage, including the injection setup. It is assumed that impurities, present in trace amounts, do not modify the velocity field of the main gas. Translational invariance along the (axial) $z$ dimension $(\frac{\partial}{\partial z}=0)$ was assumed. This assumption relies on the (axial) orientation of the distributors and the discharge optimisation performed in the previous section, which ensures the discharge angle is nearly perpendicular to the distributor (table 3). As discussed, the outgassing rate from the end-caps can be made sub-dominant relative to that from the field cage, approximately preserving the adopted symmetry. Although the region between holes cannot be well described in this approach, the distance chosen in geometries $A$-$D$ in table 3 is sufficiently small (15-50 mm) to anticipate a negligible (local, at most) impact.
@@
The hole aperture of the inlets and local pressure difference (distributor-TPC) were set to provide (simultaneously) the desired average velocity per hole and flow per unit length, based on the values in table 3. Achieving convergence in the simulations by setting the hole velocity instead of the pressure difference as an initial condition was not generally possible. The outlets were set to a fixed pressure of 10 bar. Gravity was included in the simulation, and in particular the hydrostatic pressure was automatically added at the inlets and outlets depending on their height. Outgassing was set perpendicular to the walls and to the distributors, as shown in Fig. 4. At the surface boundaries, the ‘no slip’ condition was used [55], as is customary for solid boundaries and Newtonian fluids in macro-scale flow and non-rarefied conditions. The fluid properties used were those in the default COMSOL database; see [56] pp. 765-766 for details.
Natural convection in the TPC gas has been considered in the design, although it was not explicitly included in the simulation due to its limited impact under the specific operating conditions. While temperature gradients can induce convective circulation, the flow regime in the system is primarily dominated by controlled gas injection and forced recirculation, which minimises the influence of natural convection on impurity redistribution. To assess its impact, estimates based on the Rayleigh number indicate that, given the system’s dimensions and thermal conditions, convective patterns would develop on timescales significantly longer than those observed in our CFD simulations. Moreover, the thermal stability of the system, together with the encapsulation of heat-generating components and the low thermal conductivity of the surrounding materials, significantly reduces the magnitude of the thermal gradients responsible for convection. Therefore, the isothermal approach adopted in the model remains appropriate within the scope of the present simulation.
The mesh is composed of 701 thousand linear (Lagrange P1) elements with an average skewness of 0.9 and average growth rate of 0.87 (in both metrics 1 means a perfect element and 0 a degenerated one, see [56] p. 629). The mesh is finer closer to the holes and coarser towards the centre of the field cage. To ensure convergence, the solution was checked to be stable upon global up- and down-scaling of the mesh size and variations in the time step.
@@
The turbulent-flow ‘k-$\varepsilon$’ model was selected, being adequate for simulating single-phase flows at high Reynolds numbers and suitable for incompressible or weakly compressible flows, and compressible flows at low Mach numbers. The ‘k-$\varepsilon$’ model is the most validated turbulence model and performs well in confined flows where the Reynolds shear stresses are important. This model uses the Reynolds-averaged Navier-Stokes equations (RANS), introducing two additional magnitudes: the turbulent energy ($k$) and the rate of turbulent energy dissipation ($\varepsilon$). These magnitudes have their corresponding differential equations coupling the velocity field ($\overrightarrow{v}$) with the viscous stress tensor K; the details can be found in [57]. The dimensionless coefficients, called closure coefficients, of the model are [51]:
@@
$\overrightarrow{J}=-D\overrightarrow{\nabla}c,$ (3.9)
@@
$\frac{\partial c}{\partial t}+\overrightarrow{\nabla}\cdot\overrightarrow{J}+\overrightarrow{v}\cdot\overrightarrow{\nabla}c=0,$ (3.10)
and diffusion coefficients are taken from the last column in table 2.
As a summary, the following boundary conditions and models were used:
  • Translational invariance (2D simulation).
  • TPC initial pressure: 10 bar ($20^{\circ}$).
  • Gas outlets: 10 bar (corrected by hydrostatics).
  • Gas inlets: pressure adjusted to provide the desired gas velocity at injection, as estimated from the 1D distributor model. The size of the hole (effectively a slit in a 2D simulation) is adjusted to provide the desired mass flow per unit length, as estimated from the 1D distributor model.
  • No slip at the walls.
  • Outgassing source perpendicular to the walls.
  • $k-\varepsilon$ model for the turbulence.
  • Fick’s law for the transport of impurities.
@@ 4.1 TPC filling
A priori, the displacement of air during the chamber filling can be performed at flow rates higher than those imposed by the recirculation compressors, if need be. In the ideal case, it might even be possible to evacuate the TPC, although for large volumes this cannot be taken for granted. Here we consider, for illustration, the process of TPC filling at a flow rate of 6.5 m3/h (‘typical’ flow), in the default geometry ($B$). It has been assumed, for simplicity, that the chamber is filled quickly (and uniformly) up to 9 bar of Ar and 1 bar of air (by making use of the large pressure difference at the start of injection) and afterwards the argon flow is set to the nominal one (which marks the time $t=0$ h in simulation). In the absence of convection, the triple-hole distributor proposed in geometry $B$ removes O2 concentration very efficiently from the region between distributors, and then the fresh argon gently displaces the Ar/O2 admixture upwards. The relatively low diffusion of O2 at high pressure and its lower mass favours its extraction from the uppermost outlets.
Figure 7: Evolution of the oxygen concentration upon starting the gas flow in the TPC ($\Phi=6.5$ m3/h, $P=10$ bar), for distributor geometry $B$. The initial value corresponds to a 21% O2 concentration in air at 1 bar, in a mixture of air at 1 bar with 9 bar of argon. The line shows the 0.1 ppm landmark, corresponding to a 10% charge loss in a cathode-anode electron transit.
Fig. 7 shows the concentration of oxygen integrated over the chamber as a function of time. It is worth noting that, in the limit of large diffusion (where O2 would be uniformly distributed over the chamber at all times), the time constant for O2 replacement would be directly $\tau_{{}_{\text{O}_{2}}}=V/\Phi$ (see appendix, section C.3), with $V$ being the chamber volume (100 m3). Reaching the 0.1 ppm landmark value, for instance, would take 188 h, approximately 4.5 times longer than observed in the CFD simulations (42 h).
Figure 8: Time lapse of the distribution of O2 concentration (logarithm of the concentration in mol/m3) in our reference 10 bar Ar/CF4 TPC for geometry $B$ (holes in ‘triplet’ configuration), under typical outgassing and flow rate conditions. The concentric circles mark the limits of the different mesh regions. The transient phase of the gas mixing can be observed mostly during the first 20 h. After that, the system approaches steady state and the concentration becomes homogeneous in the chamber. The images show clearly how the gas mixing is driven in these conditions by the velocity field, going from bottom to top.
@@ 4.2 Oxygen contamination
When studying the evolution of oxygen, two outgassing rates have been considered: i) $Q=10^{-8}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$ and ii) $Q=0.05\times 10^{-8}$ $\frac{\textnormal{mbar}\cdot{\textnormal{l}}}{\textnormal{cm}^{2}\cdot{\textnormal{s}}}$. Similarly, for the gas flow: i) $\Phi=6.5$ m3/h and ii) $\Phi=20$ m3/h. The outgassing rate was assumed constant and equal to its value at 10 h. This has a dual benefit: i) it provides a safe upper limit to the O2 concentration (compared to the anticipated $t^{-1/2}$ law) and ii) it allows a straightforward interpretation of when the system achieves stationary conditions. Details are compiled in Table 2.
Figure 9: Time lapse of the distribution of O2 concentration (logarithm of the concentration in mol/m3) in our reference 10 bar Ar/CF4 TPC for geometry $A$ (radial holes), under typical outgassing and flow rate conditions. After 24 h, steady state conditions have not been achieved yet.
@@
Fig. 8 shows the evolution of the O2 concentration under typical outgassing and flow rate conditions for geometry $B$ (triple-hole distributor). For comparison, Fig. 9 shows the results for geometry $A$ (single-hole distributor) when the holes are radially-oriented. This latter case displays the formation of a vortex. Therefore, while in the default geometry $B$ (hole triplet) the contaminants are uniformly mixed quickly thanks to the velocity field, in geometry $A$ (hole singlet) the uniformity is achieved through diffusion, involving considerably longer times. Single or double vortices are prevalent in single-hole distributor geometries, for the conditions studied, but are absent in the triple-hole ones. The implications of this observation can be summarised in Fig. 10. In the left panel, the time evolution of the O2 concentration is shown for three different distributor geometries (triple holes: blue; radial holes: orange; lined-up holes: green), corresponding to geometries $A$ and $B$ in table 3. The right panel shows the radial distribution of impurities (logarithmic X-scale) when all geometries have achieved steady state (500 h): it is completely homogeneous except in the outermost regions, where it can vary up to 30%. The triplet configuration reaches stationary conditions within 20 h, compared to almost 10 times that time for the 1-hole geometries. [^9] The stationary limit following eq. C.3 in appendix (section C.1) is $\frac{Q\cdot{A}}{P\cdot\Phi}=0.43$ ppm, to be compared with the simulated one of 0.31 ppm. The deviation can be attributed to i) the uncertainty in the flow estimate from the average velocity at the inlets and ii) the contribution from the last 5 cm of the chamber, that has a complex dynamics and has been removed here aiming at a simplified discussion (the last cm’s from the chamber are usually excluded from any physics analysis, as discussed earlier in text).
@@
It is difficult to assess the significance of the greater tendency to vortex formation in the single-hole distributors, and whether it will be maintained in an actual experiment when considering the exact geometry and the unavoidable convection currents stemming from temperature gradients. Notwithstanding, the triple-hole distributor is vortex-free and consistently achieves faster mixing with contaminants as well as more efficient sweeping of the outgassing from the chamber walls. This is illustrated in Fig. 11, where three different gas flows are compared under typical values of the outgassing rate (top figure: 2 m3/h, middle figure: 6.5 m3/h, bottom figure: 20 m3/h). The time evolution is displayed, up to 48 h, on the left column, while the radial distribution of contaminants is shown on the right column. Due to the low gas flow in the top figure (and correspondingly low discharge velocity), the time scale is too short to lead to vortex formation and the two distribution schemes perform similarly up to nearly the steady state value (1.4 ppm for these conditions). Even in the absence of vortices, the triple-hole distribution is more efficient at mixing due to better sweeping of the chamber walls. For larger gas flows (middle and bottom figures), the fluid dynamics of the radial-hole geometry is dominated by the formation of a central vortex, with impurities reaching the central region through diffusion alone (right figures).
Figure 12: Time evolution of the O2 contamination in our reference 10 bar Ar/CF4 TPC for ‘typical’ outgassing and flow rates, in the default geometry $B$ (holes in ‘triplet’ configuration) in red. A 1-D solution (‘infinite diffusion’ limit) is shown in dashed lines. The CFD solution for ‘optimistic’ outgassing values (blue) reflects a simple down-scaling, without altering the shape. See table 2 for typical and optimistic values. (To account for the numerical deviation from the steady-state limit, as discuss earlier in text, the outgassing in the 1-D solution has been downscaled by 28%.)
The results obtained in this section can be analysed by comparison with the one-dimensional solutions, which effectively consider an ‘infinite diffusion’ scenario (see appendix, section C.1). Fig. 12 shows how the observed time-dependence (red line) is reasonably well approximated by the 1-D solution (dashed line), implying that the achieved mixing rate is compatible with the ideal limit of perfect mixing. The blue line, obtained at 20 times less outgassing (‘optimistic’ case) aims at illustrating that the curves obtained by CFD reflect a global scaling factor for different outgassing rates, following as well the expectation from the 1-D solutions. The asymptotic value in Fig. 12 (red line) of 0.3 ppm would correspond to a 30% charge loss following the estimates in section 2.1. Considering the range of outgassing rates in table 2, an upper limit to the charge loss after 24 h can be situated in the range 1.5-30%, for the gas flow conditions in Fig. 2.1.